Mary is three years younger than three times Emily's age. The sum of their ages is 53. Find their ages.
Answer by Maths68(1474)

Let
Present age of Mary = m
Present age of Emily = e

Mary is three years younger than three times Emily's age
m=3e-3............(1)
The sum of their ages is 53
m+e=53.............(2)
Put the value of m from (1) to (2)
(3e-3)+e=53
3e-3+e=53
4e=53+3
4e=56
e=56/4
e=14
Put the value of e in (1)
m=3e-3
m=3(14)-3
m=42-3
m=39

Present age of Mary = m = 39
Present age of Emily = e = 14